OpenStudy (anonymous):

Match the following words with the correct expression. -10 more than thirty-four times x is equal to y times -31 plus -13 more than 42x

OpenStudy (anonymous):

Well, we have -10 more than something, so let's start with that: $$ -10 + \dots $$ Now that 'something' is actually thirty-four times 'x' so: $$ -10 + 34x$$ Now we're told this is equal to `y times -31` plus something, so first thing first: $$ -10 + 34x = -31y + \dots $$ What comes after the plus is told to be `-13 more than 42x`, so like the first part: $$ -10 + 34x = -31y + (-13 + 42x) $$ Which is simply: $$ - 10 + 34x = -31y - 13 + 42x $$ Notice that the phrase `y times -31 plus -13 more than 42x` is somewhat ambiguous. It could stand for \( y \cdot (-31) + (-13 + 42x)\) or maybe \( y \cdot (-31 + (-13 + 42x))\). However in other examples I've seen it means the former. If you have to pick out of given options then they probably wouldn't be both, so pick the one that exists. Hope it helps =)